Filer Manufacturing has 8,965,201 shares of common stock outstanding. The current share price is $27.09, and the book value per share is $3.78. Filer Manufacturing also has two bond issues outstanding. The first bond issue has a face value of $73,031,874, has a 0.05 coupon, matures in 10 years and sells for 83 percent of par. The second issue has a face value of $77,711,671, has a 0.06 coupon, matures in 20 years, and sells for 92 percent of par.

The most recent dividend was $0.81 and the dividend growth rate is 0.06. Assume that the overall cost of debt is the weighted average of that implied by the two outstanding debt issues. Both bonds make semiannual payments. The tax rate is 0.21.

What is Filer's WACC? Enter the answer with 4 decimals (e.g. 0.2345)
Answer #1
MV of equity=Price of equity*number of shares outstanding
MV of equity=27.09*8965201
=242867295.09
MV of Bond1=Par value*bonds outstanding*%age of par
MV of Bond1=73031874*1*0.83
=60616455.42
MV of Bond2=Par value*bonds outstanding*%age of par
MV of Bond2=77711671*1*0.92
=71494737.32
MV of firm = MV of Equity + MV of Bond1+ MV of Bond 2
=242867295.09+60616455.42+71494737.32
=374978487.83
Weight of equity = MV of Equity/MV of firm
Weight of equity = 242867295.09/374978487.83
W(E)=0.6477
Weight of debt = MV of Bond/MV of firm
Weight of debt = 132111192.74/374978487.83
W(D)=0.3523
Cost of equity
As per DDM
Price = recent dividend* (1 + growth rate )/(cost of equity - growth rate)
27.09 = 0.81 * (1+0.06) / (Cost of equity - 0.06)
Cost of equity% = 9.17
Cost of debt
Bond1
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =10x2
60616455.42 =∑ [(5*73031874/200)/(1 + YTM/200)^k]     +   73031874/(1 + YTM/200)^10x2
                   k=1
YTM1 = 7.440163436
Bond2
                  K = Nx2
Bond Price   =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                   K =20x2
71494737.32 =∑ [(6*77711671/200)/(1 + YTM/200)^k]     +   77711671/(1 + YTM/200)^20x2
                    k=1
YTM2 = 6.73
Firm cost of debt=YTM1*(MV bond1)/(MV bond1+MV bond2)+YTM2*(MV bond2)/(MV bond1+MV bond2)
Firm cost of debt=7.440163436*(60616455.42)/(60616455.42+71494737.32)+6.73*(60616455.42)/(60616455.42+71494737.32)
Firm cost of debt=7.06%
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 7.06*(1-0.21)
= 5.5774
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=5.58*0.3523+9.17*0.6477
WACC =7.91%